Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability... again

Expert replies
by Halimah_O » Mon Dec 05, 2016 6:35 am
A certain bag of gemstones is composed of two-thirds diamonds and one-third rubies. If the probability of randomly selecting two diamonds from the bag, without replacement is 5/12, what is the probability of selecting two rubies from the bag, without replacement.

(a) 5/36
(b) 5/24
(c) 1/12
(d) 1/6
(e) 1/4

IMO B[/spoiler]
Join the discussion
Source: — Problem Solving |

by regor60 » Mon Dec 05, 2016 8:24 am
Halimah_O wrote:A certain bag of gemstones is composed of two-thirds diamonds and one-third rubies. If the probability of randomly selecting two diamonds from the bag, without replacement is 5/12, what is the probability of selecting two rubies from the bag, without replacement.

(a) 5/36
(b) 5/24
(c) 1/12
(d) 1/6
(e) 1/4

IMO B[/spoiler]
The probability of selecting the first diamond is 2/3. . What you don't know is the probability of selecting the second diamond, but you're told that the probability of selecting two diamonds is 5/12.

5/12 = 2/3 multiplied by X ( probability of selecting second diamond)

X = 5/8. Now you know that because of no replacement, there could have been originally 6 diamonds (5+1) and 9 (8+1) gemstones, so 3 rubies.

With 3 rubies, probability of selecting the first one is 3/9. Probability of selecting second one is 2/8.

Probability of selecting two rubies is therefore 3/9 x 2/8 = 1/12
Join the discussion

by DavidG@VeritasPrep » Mon Dec 05, 2016 8:39 am
Halimah_O wrote:A certain bag of gemstones is composed of two-thirds diamonds and one-third rubies. If the probability of randomly selecting two diamonds from the bag, without replacement is 5/12, what is the probability of selecting two rubies from the bag, without replacement.

(a) 5/36
(b) 5/24
(c) 1/12
(d) 1/6
(e) 1/4

IMO B[/spoiler]
You could test cases until we find one that satisfies the conditions of the problem. We know that if the bag is 2/3 diamonds and 1/3 rubies that there are twice as many diamonds as rubies.

Case 1: 4 Diamonds and 2 Rubies
P(selecting 2 diamonds) = (4/6) * (3/5) = 12/30 = 2/5. We want 5/12. Close

Case 2: 6 Diamonds and 3 Rubies
P(selecting 2 diamonds) = (6/9) * (5/8) = 30/72 = 5/12. That's what we want.

If there are 3 rubies and 6 diamonds, P(selecting 2 rubies) = (3/9) * (2/8) = 6/72 = 1/12. Answer is C (Please edit the original post.)
Veritas Prep | GMAT Instructor

Veritas Prep Reviews
Save $100 off any live Veritas Prep GMAT Course
Join the discussion

by DavidG@VeritasPrep » Mon Dec 05, 2016 8:43 am
Halimah_O wrote:A certain bag of gemstones is composed of two-thirds diamonds and one-third rubies. If the probability of randomly selecting two diamonds from the bag, without replacement is 5/12, what is the probability of selecting two rubies from the bag, without replacement.

(a) 5/36
(b) 5/24
(c) 1/12
(d) 1/6
(e) 1/4

IMO B[/spoiler]
Algebraically

Diamonds: 2x
Rubies: x
Total: 3x

P(selecting 2 diamonds) = (2x/3x) * (2x-1)/(3x-1) = 5/12 --> (The x's in 2x/3x will cancel out, then we can cross multiply) --> x = 3

If x = 3,
Diamonds: 2x = 2*3 = 6
Rubies: x = 3
Total: 3x = 3*3 = 9

If there are 3 rubies and 6 diamonds, P(selecting 2 rubies) = (3/9) * (2/8) = 6/72 = 1/12. Answer is C
Veritas Prep | GMAT Instructor

Veritas Prep Reviews
Save $100 off any live Veritas Prep GMAT Course
Join the discussion

by Matt@VeritasPrep » Thu Dec 08, 2016 8:22 pm
If we know that D * (D - 1) / ((D + R) * (D + R - 1)) = 5/12, we could look for friendly values of D that give us this result.

Start by looking for two consecutive integers whose product is a multiple of 12. (This is what we need for the denominator.) 3 * 4 works, but then the numerators won't work out. (We can't get integer * integer = 5 if both of those integers are less than 4!)

The next possibility is 8 * 9. If our answer is x/72 = 5/12, then x = 30. We need D * (D - 1) = 30, so D = 6, and then, voila, we're set!

D = 6, and (D + R) = 9, so there are six diamonds and three rubies.

From there, the probability of selecting two rubies = (3 / 9) * (2 / 8) = 1/12.
Join the discussion

by Matt@VeritasPrep » Thu Dec 08, 2016 8:25 pm
Another approach:

We know there are half as many rubies as diamonds, so R = D/2.

From there, we're told that

D/(D + R) * (D - 1)/(D + R - 1) = 5/12.

Substituting D for 2R, we get

2R/(2R + R) * (2R - 1)/(2R + R - 1) = 5/12

or

2/3 * (2R - 1)/(3R - 1) = 5/12

or

(2R - 1)/(3R - 1) = 15/24 = 5/8

and

R = 3

From there, we know R = 3, D = 6, and the computation is easy.
Join the discussion

by Anaira Mitch » Thu Jan 05, 2017 7:27 pm
The simplest way to solve the problem is to recognize that the total number of gems in the bag must be a multiple of 3, since we have 2/3 diamonds and 1/3 rubies. If we had a total number that was not divisible by 3, we would not be able to divide the stones into thirds. Given this fact, we can test some multiples of 3 to see whether any fit the description in the question.
The smallest number of gems we could have is 6: 4 diamonds and 2 rubies (since we need at least 2 rubies). Is the probability of selecting two of these diamonds equal to 5/12?
4/6 × 3/5 = 12/30 = 2/5. Since this does not equal 5/12, this cannot be the total number of gems.
The next multiple of 3 is 9, which yields 6 diamonds and 3 rubies:
6/9 × 5/8 = 30/72 = 5/12. Since this matches the probability in the question, we know we have 6 diamonds and 3 rubies. Now we can figure out the probability of selecting two rubies:
3/9 × 2/8 = 6/72 = 1/12
The correct answer is C.
Join the discussion

by [email protected] » Fri Jan 06, 2017 10:30 am
Hi All,

We can solve this problem by TESTing VALUES. However, we have so much specific information, we CANNOT TEST random values. We have to use the information in the prompt to pick a logical number that matches all of the given "restrictions"

Here's what we have to work with:
1) Since the gems can be broken down into 2/3 diamonds and 1/3 rubies, the TOTAL must be a MULTIPLE of 3.
2) Since the probability of pulling 2 diamonds is 5/12, when we multiply the two individual probabilities, we MUST end with a denominator that is a multiple of 12 (so the fraction can be reduced to 5/12).

Let's start at "3" and work up....

If there are 3 gems, then we have 2 diamonds.
The probability of pulling 2 diamonds is (2/3)(1/2) = 2/6 which is NOT a match.

If there are 6 gems, then we have 4 diamonds.
The probability of pulling 2 diamonds is (4/6)(3/5) = 12/30.....30 cannot reduce to 12. This is NOT a match

If there are 9 gems, then we have 6 diamonds.
The probability of pulling 2 diamonds is (6/9)(5/8) = 5/12...This IS a MATCH

So we have....
Total= 9
Diamonds = 6
Rubies = 3

The question asks for the probability of selecting 2 rubies....

The probability of selecting the first ruby = (3/9)
The probability of selecting the second ruby = (2/8)
(3/9)(2/8) = 6/72 = 1/12

Final Answer: C

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion